Q1: Find the general solution to the ODE:

y'' + 4y' + 13y = 0.
Answer #1

The given ordinary differential equation is

y'' + 4y' + 13y = 0

CD?+40 + 13)0 where D=d da The aurilary equation is m2 + 4m +13°=0 → m = -4+42_4x1x13 - 2X1 3-4 + J16-52 - 2 =-4+61 2 m = -2+

which is the required general solution of given ordinary differential equation.